OpenStudy (anonymous):

what is the solution of the system of equations? x+4y=7 -3x+y=-8

OpenStudy (anonymous):

y = -5 x = 27

OpenStudy (anonymous):

sorry i messed up the equation @mhescott

OpenStudy (anonymous):

and can u explain how u got the answer

OpenStudy (anonymous):

solve for either x or y. then plug that answer into the other equation. using the first equation: x=-4y+7. Now use that in the other equation by plugging this equation where x is. doing that you'll solve that equation for y coming up with y=1. Now plug y into the first equation and solve for x. doing so you'll find that x=3 Always check the answer by plugging the answers youve gotten back into either equation.

OpenStudy (anonymous):

I don't agree with mhescott. What you want to do is to try to get rid of one of the terms in one of the equations, so that you're only working with one variable. You can see that if you multiplied the first equation by 3, you'd have 3x + 12y = 21. Then you could add both equations (3x + 12y = 21 and -3x + y = -8) together. The 3x and -3x would =0. the 12y + y = 13y and 21 -8 = 13. You are left with 13y = 13. Multiply both sides by 1/13 and you get y = 1. Now substitute 1 in each equation for y and you'll get x. You'll see that the x in both equations is the same. There are other approaches to use, but this one is the easiest. Hope this helps.